OpenStudy (anonymous):

The x-intercept of 3x + 2y = 12 is (__, 0). so im trying to find x right????

OpenStudy (anonymous):

put y=0 3x=12 x= 12/3=4

OpenStudy (anonymous):

3(12)+2y=12??

OpenStudy (anonymous):

3(4)+2*(0)=12

OpenStudy (anonymous):

oh ok but why is x 4?

OpenStudy (anonymous):

put y=0 you will get 3x= 12 now divide both sides by 3 \[\frac{ 3x }{ 3 }=\frac{ 12 }{ 3 }=4=x\]

OpenStudy (anonymous):

alright thanks! :)

OpenStudy (anonymous):

so the answer to The y-intercept for 3x + 2y = 12 is (0,___). would be 6

OpenStudy (anonymous):

yep (:

OpenStudy (anonymous):

thanks :)